# TigerLLM - A Family of Bangla Large Language Models

### v2

@@ Abstract
Nishat Raihan, Marcos Zampieri George Mason University, VA, USA mraihan2@gmu.edu
@@ 1 Introduction
LLMs have fundamentally transformed NLP by achieving exceptional performance across a broad range of tasks Brown et al. (2020); Raihan et al. (2025); Zhang et al. (2023). While these models exhibit unprecedented capabilities in language understanding, generation, reasoning, and specialized domains Raihan et al. (2024b), their advancements predominantly benefit high-resource languages Alam et al. (2024). This inequality is particularly noticeable for Bangla. Despite having about 237 million native speakers,111ethnologue.com/language/ben/ Bangla remains quite underserved in modern NLP advancements.
This under-representation stems primarily from the limitation of high-quality training data. While proprietary models like GPT-4 Brown et al. (2023) and Claude-3.5 Bai et al. (2024) demonstrate reasonable Bangla capabilities, open-source alternatives consistently underperform. Recent multilingual models such as Gemma-2 Team et al. (2024) and LLaMA 3.1 Dubey et al. (2024), despite leveraging diverse training corpora and advanced tokenization systems like TikTokenizer Corso et al. (2024), also fail to deliver satisfactory performance for Bangla.
@@ 1.1 Limitations of Bangla LLM Initiatives
Recent attempts at developing Bangla LLMs (see Table 1) through continual pretraining titu-Gemma222huggingface.co/hishab/titulm-gemma-2-2b-v1.1 and model distillation approaches Zehady et al. (2024) have yielded low and non-reproducible results (see Table 2), often performing worse than their base models. The absence of technical documentation and academic publications further compounds this issue by making result reproduction impossible. Our investigation into these models’ performances reveals the need for improvement in the training process. While the unavailability of pretraining corpora limits our analysis of that phase, the finetuning approach demonstrates consistent problematic patterns.
@@
Most Bangla LLM initiatives rely on translated versions of synthetic datasets like Alpaca-Instruct Taori et al. (2023) and OpenOrca Mitra et al. (2023), which are generated through model distillation Hinton et al. (2015). This approach suffers from two fundamental limitations: (1) the datasets are generated by early GPT-3.5 Brown et al. (2020) releases, a model with limited Bangla support, resulting in suboptimal instruction quality, and (2) these English datasets are translated to Bangla using machine translation systems like Google Translate 333translate.google.com with limited quality checks, further degrading the training data quality. These cascading compromises in training data ultimately result in poor model performance.
@@ 1.2 Contributions
All components are open-sourced to establish robust foundations for future Bangla language modeling research.444github.com/mraihan-gmu/TigerLLM/tree/main
@@ 2 Bangla-TextBook Corpus
Previous Bangla LLMs rely predominantly on corpora sourced from OSCAR Ortiz Suárez et al. (2020) and Common Crawl Bhattacharjee et al. (2022); Zehady et al. (2024), despite quality control challenges. While alternative Bangla corpora have emerged Bhattacharyya et al. (2023); Sketch Engine (2021), the absence of curated educational content remains a critical gap. This emphasis on data quality is particularly significant given recent findings by Gunasekar et al. (2023), which demonstrate that LLMs achieve superior performance through high-quality training data, even with reduced volume.
To bridge this gap, we present the Bangla-TextBook corpus, constructed exclusively from high-quality open-source educational materials published by the National Curriculum and Textbook Board of Bangladesh. We collect texts from 163 textbooks for Grades 6-12 resulting in a total of a total of 9,897,623 tokens and 697,903 sentences.
@@ 3 Bangla-Instruct
1.
New Instruction Generation: Let be our seed task pool (Appendix A.2. For each generation round , we sample tasks uniformly at random from as demonstrations. Claude generates new instructions conditioned on these demonstrations while preserving Bangla linguistic patterns.
2.
Task Identification: Each new instruction undergoes classification by GPT-4o to determine task type . This classification helps in generating appropriate response formats.
3.
Instruction-Response Generation: For each instruction , we generate comprehensive responses using Claude that generates a candidate response.
4.
Filtering: Finally, GPT-4o carries out the multi-stage filtering process ensuring the Bengali language adherence, Cultural sensitivity, Response completeness and Discarding Similar instances (see Appendix A.3).
Valid instruction-response pairs are added back to the task pool, enabling iterative refinement. This process continues until we reach our target dataset size of 100K pairs, with each pair satisfying our quality criteria. The dual-model approach, combined with rigorous filtering, significantly reduces the likelihood of generating inappropriate or low-quality content, a common challenge in previous Bengali instruction datasets. More details are presented in Appendix A.
4 TigerLLM
We consider multiple candidate base models, similar to Raihan et al. (2024c)’s approach. 3 families of multilingual LLMs are chosen - LLaMA 3.2 (1B, 3B) Dubey et al. (2024), Gemma-2 (2B, 9B) Team et al. (2024) and Pangea (7B) Yue et al. (2024). Figure 2 depicts the final selection of model and a high-level overview of the process.
Continual Pretraining
We use the Bangla-TextBook corpus for the models to learn culture and language-specific nuances and gather sufficient and reliable knowledge from a set of high-quality texts. The pretraining phase has been carried out multiple times with empirical choices of hyper-parameters. the final set of hyperparameters is presented in Appendix B.
Model Distillation
We finetune our pretrained models using Bangla-Instruct. LoRA Hu et al. is not used, we implement full finetuning for better learning. To speed up the training process, we utilize Flash Attention Dao et al. (2022), we set key parameters: 2048 token maximum sequence length, batch size of 8, 4 gradient accumulation steps, and 3 epochs. Learning rate (), weight decay (0.02), and 10% warm-up steps ensure stable convergence. Table 5 in Appendix B lists complete hyperparameters.
@@ 5 Evaluation
Although there has been limited research on Bangla LLMs, several benchmarks have been established to assess their performance. We focus on five benchmarks specifically curated to evaluate Bangla LLMs across a diverse set tasks. For multitask understanding, we use the Bangla subset of MMLU-Pro Wang et al. (2024) and PangBench Yue et al. (2024). For question answering, we consider BanglaQuaD Rony et al. (2024), while for general knowledge, we use BEnQA Shafayat et al. (2024). For reasoning tasks, we refer to BanglaRQA Ekram et al. (2022), and for coding tasks, we utilize the Bangla subset of mHumanEval Raihan et al. (2024a).
@@ 6 Conclusion and Future Work
This paper introduces TigerLLM, a family of state-of-the-art Bangla language models that outperforms existing alternatives across six benchmarks. TigerLLM’s success stems from two key innovations: (1) the high-quality Bangla-TextBook corpus derived from educational materials, and (2) the carefully curated Bangla-Instruct dataset generated using advanced teacher models. The three resources introduced here (corpus, instruct dataset, and models) establish a robust foundation for future Bangla language modeling research. Together they will contribute to speed up advances in Bangla language modeling.
Future work will focus on (1) qualitative analysis of model performance, (2) expanding the corpus to include more diverse domains, (3) scaling up model sizes while maintaining quality, and (4) developing more sophisticated evaluation metrics specific to Bangla language tasks.
@@
While TigerLLM demonstrates impressive performance, several limitations warrant acknowledgment. First, our Bangla-TextBook corpus, though carefully curated, is limited to educational materials from grades 6-12, potentially missing broader linguistic patterns present in other domains. The 10 million token size, while sufficient for our current models, may constrain scaling to larger architectures. Additionally, our Bangla-Instruct dataset, despite its quality-focused generation process, covers only a subset of possible instruction types and may not fully capture the complexity of real-world Bangla language use cases.
@@ Appendix B Experimentation Details
B.1 Experimental Setup
Pretraining
We utilize a Lambda Labs555lambdalabs.com cluster with 8 NVIDIA A100 GPUs (40GB each), 512GB RAM, and 2TB storage. The distributed training setup enables efficient parallel processing, completing the pretraining in approximately 120 hours on this high-performance configuration with gradient checkpointing enabled.
Finetuning
We conduct finetuning on a single NVIDIA A100 (40GB) through Google Colab666colab.research.google.com, supported by 80GB RAM and 256GB storage. The process completes in approximately 96 hours, proving sufficient for model adaptation and task-specific optimization with minimal computational overhead.
B.2 Pretraining HyperParameters
@@
B.3 Finetuning Hyperparameters
| Hyperparameter | Value |
@@ B.3 Finetuning Hyperparameters
Appendix C TigerLLM - Training Pipeline
Figure 2 illustrates the multi-stage training pipeline for producing both TigerLLM (1B) and TigerLLM (9B). It begins with two pre-trained language models—LLaMA 3.2 (1B) and Gemma 2 (9B)—chosen for their robust foundational capacities. These models then undergo continual pretraining (see Figure 3) on a specialized Bangla-TextBook corpus, which infuses them with a richer understanding of the Bangla language, including its context-specific nuances, stylistic variations, and domain-specific terminology.
Following this continual pretraining step (Figure 3), the models are finetuned on a carefully curated Bangla-Instruct dataset. This additional phase (Figure 4) ensures that the trained models can effectively process and respond to Bangla instructions, yielding outputs that are more coherent, context-aware, and aligned with users’ informational needs.
In essence, by blending the foundational strengths of LLaMA and Gemma with specialized Bangla corpora and instruction-oriented finetuning, the final TigerLLM models emerge as optimized solutions capable of delivering high-quality, instruction-following responses tailored to Bangla-language tasks.

### v3

@@ Abstract
Nishat Raihan George Mason University Fairfax, VA, USA mraihan2@gmu.edu Marcos Zampieri George Mason University Fairfax, VA, USA mzampier@gmu.edu
@@ 1 Introduction
LLMs have fundamentally transformed NLP by achieving exceptional performance across a broad range of tasks Brown et al. (2020); Chowdhery et al. (2022); Raihan et al. (2025c). While these models exhibit unprecedented capabilities in language understanding, generation, reasoning, and specialized applications, their advancements predominantly benefit high-resource languages Alam et al. (2024). This inequality is particularly noticeable for Bangla. Despite having about 237 million native speakers,111ethnologue.com/language/ben/ Bangla remains quite underserved in modern NLP advancements.
This under-representation stems primarily from the limitation of high-quality training data. While proprietary models like GPT-4 Brown et al. (2023) and Claude-3.5 Bai et al. (2024) demonstrate reasonable Bangla capabilities, open-source alternatives consistently underperform. Recent multilingual models such as Gemma-2 Gemma et al. (2024) and LLaMA 3.1 Dubey et al. (2024), despite leveraging diverse training corpora and advanced tokenization systems like TikTokenizer Corso et al. (2024), also fail to deliver satisfactory performance for Bangla.
@@ 1.1 Limitations of Bangla LLM Initiatives
Recent attempts at developing Bangla LLMs (see Table 1) through continual pretraining (titu-Gemma) and model distillation approaches Zehady et al. (2024) have yielded low and non-reproducible results (see Table 2), often performing worse than their base models. The absence of technical documentation and academic publications further compounds this issue by making result reproduction impossible. Our investigation into these models’ performances reveals the need for improvement in the training process. While the unavailability of pretraining corpora limits our analysis of that phase, the finetuning approach demonstrates consistent problematic patterns.
@@
Most Bangla LLM initiatives rely on translated versions of synthetic datasets like Alpaca-Instruct Taori et al. (2023) and OpenOrca Mitra et al. (2023), which are generated through model distillation Hinton et al. (2015). This approach suffers from two fundamental limitations: (1) the datasets are generated by early GPT-3.5 Brown et al. (2020) releases, a model with limited Bangla support, resulting in suboptimal instruction quality, and (2) these English datasets are translated to Bangla using machine translation systems like Google Translate with limited quality checks, further degrading the training data quality. These cascading compromises in training data ultimately result in poor model performance.
@@ 1.2 Contributions
All components are open-sourced to establish robust foundations for future Bangla language modeling research.222https://github.com/mraihan-gmu/TigerLLM/tree/main/
2 Related Work
Early transformer-based encoder-only pre-trained language models such as BERT Devlin et al. (2019) concentrate on high-resource languages like English. Subsequent work adapts them to mid- and low-resource contexts through continued pre-training and task-specific finetuning. In Bangla, for instance, Sami et al. (2022) present BanglaBERT, demonstrating that a dedicated monolingual encoder markedly improves downstream classification and QA relative to multilingual baselines.
The shift to decoder-only models has produced large multilingual models — e.g. BLOOM Le Scao et al. (2022), Llama3 Dubey et al. (2024), and Aya Üstün et al. (2024)—that cover dozens of under-represented languages. Yet empirical analyses reveal that these models still perform best when prompted in high-resource languages, with significant degradation for languages such as Bangla or Swahili Raihan et al. (2025a); Jin et al. (2024).
As discussed in the previous section, dedicated Bangla decoder models remain scarce and fragmented. GPT2-Bangla Bhattacharjee et al. (2023) continues GPT-2 pre-training on a 4GB Bangla corpus, while Bong-LLaMA Zehady et al. (2024) and the titu-Gemma333https://huggingface.co/hishab/titulm-gemma-2-2b-v1.1 checkpoint attempt instruction tuning on translated datasets. These efforts often lack rigorous evaluation protocols, transparent data curation, or reproducible training pipelines—as reflected in the inconsistent results summarized in Table 1. Consequently, a clear methodological gap persists in developing open, reproducible decoder-only LLMs that natively support Bangla and other low-resource languages.
@@ 3 Bangla-TextBook Corpus
Previous Bangla LLMs rely predominantly on corpora sourced from OSCAR Ortiz Suárez et al. and Common Crawl Bhattacharjee et al. (2022); Zehady et al. (2024), despite quality control challenges. While alternative Bangla corpora have emerged Bhattacharyya et al. (2023), the absence of curated educational content remains a critical gap. This emphasis on data quality is particularly significant given recent findings by Gunasekar et al. (2023) and Raihan et al. (2025b), which demonstrate that LLMs achieve superior performance through high-quality training data, even with reduced volume.
To bridge this gap, we present the Bangla-TextBook corpus, constructed exclusively from high-quality open-source educational materials published by the National Curriculum and Textbook Board of Bangladesh. We collect texts from 163 textbooks for Grades 6-12, resulting in a total of 9,897,623 tokens and 697,903 sentences.
@@ 4 Bangla-Instruct
(1) Seed & Instruction Generation: We begin with a human-curated seed pool drawn from 50 volunteers representing five academic disciplines across Bangladesh (see Appendix A.1). At every generation round , we sample seed tasks and prompt Claude to create a candidate batch of instructions , expanding coverage of the ten seed categories listed in Appendix A.2 while preserving authentic linguistic patterns.
(2) Task Typing: Each instruction is classified by GPT-4o into , providing the expected answer style and the minimum-length threshold used in subsequent filtering.
(3) Response Drafting: Conditioned on , Claude produces a comprehensive response . We retain the highest-scoring draft according to an internal coherence metric .
(4) Multi-stage Filtering: GPT-4o applies the four-criteria filter —Language (), Cultural (), Quality (), and Novelty () (see Appendix A.3). On average, ~63% of pairs pass , yielding a balanced complexity mix (40% basic, 40% intermediate, 20% advanced). Valid pairs are appended to , and the loop continues until 100K high-quality instruction–response pairs are reached.
By coupling two complementary LLMs with strict verification and a human-seeded, domain-balanced task pool, our pipeline mitigates error propagation and preserves cultural nuance—addressing shortcomings observed in earlier Bengali instruction datasets (see Appendix A for full statistics).
5 TigerLLM
As candidate base models, we consider 3 families of multilingual LLMs - LLaMA 3.2 (1B, 3B) Dubey et al. (2024), Gemma-2 (2B, 9B) Gemma et al. (2024) and Pangea (7B) Yue et al. (2024).
Evolution of TigerLLM
Figure 2 depicts the final selection of the models and a high-level overview of the process.
Upon the selection phase, we finalize two pre-trained language models—LLaMA 3.2 (1B) and Gemma 2 (9B)—chosen for their robust foundational capacities. These models then undergo continual pretraining (see Figure 3) on a specialized Bangla-TextBook corpus, which infuses them with a richer understanding of the Bangla language, including its context-specific nuances, stylistic variations, and domain-specific terminology.
Pretraining
We utilize a computing cluster with 8 NVIDIA A100 GPUs (40GB each), 512GB RAM, and 2TB storage. The distributed training setup enables efficient parallel processing, completing the pretraining in approximately 120 hours on this high-performance configuration with gradient checkpointing enabled.
Continual Pretraining
We use the Bangla-TextBook corpus for the models to learn culture and language-specific nuances and gather sufficient and reliable knowledge from a set of high-quality texts. The pretraining phase has been carried out multiple times with empirical choices of hyper-parameters.
Finetuning
We conduct finetuning on a single NVIDIA A100 (40GB) through Google Colab444colab.research.google.com, supported by 80GB RAM and 256GB storage. The process completes in approximately 96 hours, proving sufficient for model adaptation and task-specific optimization with minimal computational overhead.
Model Distillation
Following this continual pretraining step, the models are finetuned on a carefully curated Bangla-Instruct dataset (Figure 4). LoRA Hu et al. (2021) is not used, we implement full finetuning for better learning. To speed up the training process, we utilize Flash Attention Dao et al. (2022), we set key parameters: 2048 token maximum sequence length, batch size of 8, 4 gradient accumulation steps, and 3 epochs. Learning rate (), weight decay (0.02), and 10% warm-up steps ensure stable convergence. Table 5 in Appendix B lists complete hyperparameters.
@@ 5 TigerLLM
By blending the foundational strengths of LLaMA and Gemma with specialized Bangla corpora and instruction-oriented finetuning, the final TigerLLM models emerge as optimized solutions capable of delivering high-quality, instruction-following responses tailored to Bangla-language tasks.
@@ 6 Evaluation
Although there has been limited research on Bangla LLMs, several benchmarks have been established to assess their performance. We focus on five benchmarks specifically curated to evaluate Bangla LLMs across a diverse set tasks. For multitask understanding, we use the Bangla subset of MMLU-Pro Wang et al. (2024) and PangBench Yue et al. (2024). For question answering, we consider BanglaQuaD Rony et al. (2024), while for general knowledge, we use BEnQA Shafayat et al. (2024). For reasoning tasks, we refer to BanglaRQA Ekram et al. (2022).
@@ 7 Conclusion and Future Work
This paper introduces TigerLLM, a family of state-of-the-art Bangla language models that outperforms existing alternatives across six benchmarks. TigerLLM’s success stems from two key innovations: (1) the high-quality Bangla-TextBook corpus derived from educational materials and (2) the carefully curated Bangla-Instruct dataset generated using advanced teacher models.
The three resources introduced here (corpus, instruction dataset, and models) establish a robust foundation for future Bangla language modeling research. Together, they will contribute to speeding up advances in Bangla language modeling.
In future work we will conduct a deeper qualitative analysis of the model’s behavior, broaden the corpus to cover a wider array of domains, scale the model to larger parameter counts without compromising quality, and devise richer evaluation metrics tailored specifically to Bangla tasks.
@@
While TigerLLM delivers state-of-the-art performance, several limitations warrant acknowledgment. First, our Bangla-TextBook corpus, though carefully curated, is limited to educational materials from grades 6-12, potentially missing broader linguistic patterns present in other domains. The 10 million token size, while sufficient for our current models, may constrain scaling to larger architectures. Additionally, our Bangla-Instruct dataset, despite its quality-focused generation process, covers only a subset of possible instruction types and may not fully capture the complexity of real-world Bangla language use cases.
@@ Appendix B Experimentation Details
B.1 Pretraining HyperParameters
@@
B.2 Finetuning Hyperparameters
| Parameter | Value |